### v4

# Task Aware Dreamer for Task Generalization in Reinforcement Learning

Abstract
A long-standing goal of reinforcement learning is to acquire agents that can learn on various training tasks and generalize well on unseen tasks that may share a similar dynamic but with different reward functions. The ability to generalize across tasks is significant for real-world applications where the robot needs to adapt to varying reward mechanisms with the same embodiment. In this work, we first indicate that training general world models can utilize similar structures in these tasks and benefit training more generalizable agents. Extending world models into the task generalization setting, we introduce a novel method named Task Aware Dreamer (TAD), which integrates reward-informed features to identify consistent latent characteristics across tasks. Within TAD, we compute the variational lower bound of sample data log-likelihood, which introduces a new term designed to differentiate tasks using their states, as the optimization objective of our reward-informed world models. To demonstrate the advantages of the reward-informed policy utilized in TAD for handling the task distribution, we introduce a novel metric named Task Distribution Relevance (TDR) which quantitatively measures the relevance of different tasks. For tasks exhibiting a high TDR, i.e., these tasks differ significantly, we illustrate that Markovian policies struggle to distinguish them, thus it is necessary to utilize reward-informed policies in TAD. Extensive experiments in both image-based and state-based settings show that TAD can significantly improve the performance of handling different tasks meanwhile, especially for those with high TDR, and display a strong generalization ability to unseen tasks.
Figure 1: An overview. Given a task distribution, we train the agent in training tasks and hope it to zero-shot generalize to test tasks. For improving the generalization, we propose TAD, which utilizes $\Pi_{3}$ to encode all historical information for inferring the current task and novel reward-informed world models for capturing invariant latent features.
1 Introduction
Deep Reinforcement Learning (DRL) has demonstrated significant advancements in diverse fields (Mnih et al. 2016; Silver et al. 2016), and a key factor in these achievements is an agent’s proficiency in assimilating lessons from special training tasks. This tendency towards specialization harms the wider real-world application of DRL, where broad generalization across various tasks is necessary. One primary challenge in generalization is the correspondence in the underlying dynamics of different tasks, which exhibit distinct reward structures. For example, we may need to control the same robot to handle different tasks, which can be rewarded for speed or caution in a navigating robot. Developing generalizable agents that recognize and handle these subtle variations is still an area of keen interest and exploration in DRL.
For developing such generalizable agents, a promising pathway is to train general world models (Ha and Schmidhuber 2018; Hafner et al. 2019b, a) that help the agent understand the world and make decisions. In this work, we first provide insight that general world models benefit improving the sample efficiency in handling the task distribution, especially by utilizing similar dynamic structures of trajectories sampled from different tasks (Theorem 1). Consequently, we propose a novel framework named Task Aware Dreamer (TAD) to improve the generalization ability of agents via general world models. As current world models are primarily designed for the single-task setting (left of Fig. 2), we consider the corresponding probabilistic graphical model for the task-distribution setting (right of Fig. 2) and propose novel reward-informed world models that can capture invariant structures across tasks in the distribution. Then we compute the variational lower bound of the data log-likelihood as the primary training objective for optimizing the reward-informed world models, which incorporate a novel task context term that categorizes tasks based on their states. In practice, we implement this optimization objective via two alternative methods: cross-entropy (TAD-CE) and supervised-contrastive (TAD-SC).
Figure 2: Probabilistic graphical model designs for the single-task setting (left) and the task-distribution setting (right). The latter inspires the design of reward-informed world models. Solid and dashed lines represent the generative process and the inference model, respectively.
Moreover, we theoretically explain TAD’s components, like reward-informed policies and task optimization terms, are effective for handling task distribution. About the policy hypothesis, although several meta RL methods (Zintgraf et al. 2019; Rimon et al. 2024) have chosen the reward-informed policy hypothesis $\Pi_{3}$, the relationship between the task distribution and the expressiveness of the policy hypothesis set is still unclear. In other words, why the commonly adopted hypothesis set $\Pi_{1}$ of Markovian policies and the set $\Pi_{2}$ of policies encoding historical states and actions are not suitable for task distribution? To answer this question, we propose a novel metric of Task Distribution Relevance (TDR), encapsulating the relevance of different tasks within the distribution through their optimal Q functions. We then prove that both $\Pi_{1}$ and $\Pi_{2}$ are sub-optimal under the task-distribution setting. This sub-optimality is related to TDR (Theorem 3), i.e., for task distributions with high TDR, the performance of these two policy hypotheses might degenerate significantly, a phenomenon we also demonstrate in experiments (Sec. 5.2). This result explains why TAD and previous meta RL methods choose $\Pi_{3}$ for handling the task distribution. Besides the policy chosen, we also discuss our task optimization term in TAD and prove that it can effectively reduce the gap between the policy return and the optimal return (Theorem 4).
We evaluate the task generalization ability of TAD in extensive experiments, including DeepMind control suite (Tassa et al. 2018) and MuJoCo (Todorov, Erez, and Tassa 2012), which are image-based and state-based respectively. Agents are trained on various tasks and evaluated on unseen tasks. Results corroborate our analyses, indicating that $\Pi_{1}$ and $\Pi_{2}$ falter in managing task distributions characterized by a high TDR. Contrastingly, TAD excels in simultaneously managing varied tasks and outperforms all baselines, including SOTA model-based meta RL method MAMBA (Rimon et al. 2024). Additionally, our ablation studies highlight TAD’s versatility, showing its prowess in dynamic generalization and handling cross-embodiment tasks. Overall, our contributions include:
  • We present theoretical insights that general world models can utilize similar structures across tasks and improve the sample efficiency for task generalization (Sec. 4.1).
  • Our TAD extends world models for the task distribution with the corresponding variational lower bound (Sec. 4.2).
  • We theoretically analyze TAD’s components are effective for task generalization, with a novel metric TDR to quantify the distribution relevance (Sec. 4.3).
  • Extensive experiments show that TAD can outperform various SOTA baselines and exhibit better generalization flexibility over image-based and state-based settings (Sec. 5).
2 Related Work
Generalization in RL.
Current RL methods always struggle to generalize to new tasks (Song et al. 2019). Prior works have studied an array of training strategies, like loss regularization (Cobbe et al. 2019; Wang et al. 2020), successor representation (Touati and Ollivier 2021; Touati, Rapin, and Ollivier 2022), network architecture design (Lee et al. 2019; Raileanu and Fergus 2021), lifelong learning (Chen and Liu 2018; Mendez, van Seijen, and Eaton 2022), data augmentation (Raileanu et al. 2021; Hansen and Wang 2021), etc. Besides these, some works investigate the connection between policy generalization and the distribution of those environments. Ghosh et al. demonstrates that generalizing to unseen environments introduces partial observability, thereby rendering deterministic Markovian policies sub-optimal. Also, some studies (Lee et al. 2020; Ghosh et al. 2021) experimentally indicate that stochastic or non-Markovian policies can improve the generalization ability. However, the expressive abilities of differing hypothesis sets and their connection to the environment distribution, which are significant for developing more generalizable agents, remain understudied.
Multi-task RL and Meta RL.
These two topics are closely related to generalization in RL. Multi-task RL (Yang et al. 2020; Sodhani, Zhang, and Pineau 2021; Lee et al. 2022; Xu et al. 2022a) primarily aims to excel across all training tasks but is difficult to zero-shot generalize to unseen tasks. For boosting the generalization, Meta RL seeks to enable the trained agents to adapt to new tasks with few episodes, including gradient-based (Finn, Abbeel, and Levine 2017) and context-based methods (Duan et al. 2016; Rakelly et al. 2019). There are also some model-based methods (Nagabandi et al. 2018; Rimon et al. 2024) utilize learned models to boost the sample efficiency. Though some context-based methods like VariBAD (Zintgraf et al. 2019) show zero-shot generalization ability, it is still significant to directly analyze the generalization in RL and design corresponding algorithms.
World Models.
World models (Ha and Schmidhuber 2018) aims to better learn environmental representations, which has potential advantages for generalization as it can capture invariant features across tasks. Classical methods utilize the Recurrent State Space Model (RSSM) (Hafner et al. 2019b) for planning (Hafner et al. 2019b) and policy learning (Hafner et al. 2019a, 2020, 2023). Subsequent research explored reconstruction-free world models (Deng, Jang, and Ahn 2022), temporal predictive coding (Nguyen et al. 2021), cooperative reconstruction (Fu et al. 2021), and Denoised MDP (Wang et al. 2022) for more effective task-relevant information encoding. World models are also utilized to extract environmental invariant features by learning from videos (Seo et al. 2022) or exploration (Sekar et al. 2020; Xu et al. 2022b), and then fine-tuning to new tasks. However, most world models are designed for the single-task setting and struggle to manage multiple tasks without fine-tuning, limiting their effectiveness for zero-shot generalization to unseen tasks.
3 Preliminary
We consider the setting with a task distribution $\mathcal{T}$ of Partially observable Markov decision processes (POMDPs), where different tasks own the same dynamic and different rewards. Formally, each POMDP $\mathcal{M}\sim\mathcal{T}$ can be represented as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\Omega,\mathcal{O})$. Here $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. For $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathcal{P}(\cdot|s,a)$ is the state transition probability that is Markovian, $\mathcal{R}_{\mathcal{M}}(s,a)$ is its reward function. The state is unobserved for the agent, which can only access the observation from the observation space $\Omega$ calculated by the observation function $\mathcal{O}(\cdot|s)$.
Following previous meta RL and generalization methods, we consider policies that encoder all historical information (we prove its necessity in Sec. 4.3). Formally, at each timestep $t$, the agent with the policy $\pi$ will use the whole history trajectory $(o_{0},a_{0},r_{0},o_{1},...,o_{t})$ to sample action $a_{t}$, arrive at the next state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$, and get the current reward $r_{t}=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$. The performance of policy $\pi$ in $\mathcal{M}$ is defined as the expected discounted return: $J_{\mathcal{M}}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\triangleq\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$. Our objective is to maximize the return over $\mathcal{T}$, i.e., $\max_{\pi}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]$.
In practice, given the task distribution $\mathcal{T}$, we sample $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ for optimizing the agent, i.e., maximizing $\frac{1}{M}\sum_{m=1}^{M}J_{\mathcal{M}_{m}}(\pi)$. In the testing stage, we will sample $N$ unseen test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$ to evaluate its generalization ability, i.e., evaluating $\frac{1}{N}\sum_{n=1}^{N}J_{\mathcal{M}_{M+n}}(\pi)$.
4 Task Aware Dreamer
In this section, we first demonstrate that general world models benefit task generalization. Then we introduce our reward-informed world models and propose a novel framework of Task Aware Dreamer (TAD) for handling task generalization. Finally, we provide theoretical analyses of designs in TAD.
4.1 Reward-Informed World Models
Our first observation is that general world models are effective in narrowing down the hypothesis space of the optimal Q function when handling task generalization:
Theorem 1 (Proof in Appendix A.1).
Set $\mathcal{Q}$ as the space of observation-action Q functions. Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ and corresponding dataset $\mathcal{D}_{m}=\{(o_{t}^{m},a_{t}^{m},r_{t}^{m},o_{t+1}^{m})\}$, we set the product space $\mathcal{H}=\mathcal{Q}^{M}$ composed of M spaces, i.e., $\forall\{q_{m}\}_{m=1}^{M}\in\mathcal{H}$, $q_{m}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ belongs to $\mathcal{Q}$. Considering the following three hypothesis classes $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}\subseteq\mathcal{H}$:
$\begin{split}\mathcal{H}_{1}=&\{(q_{m})_{m=1}^{M}|q_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime})\}\\ \mathcal{H}_{2}=&\{(q_{m})_{m=1}^{M}|\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\}\\ \mathcal{H}_{3}=&\{(q_{m})_{m=1}^{M}|\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall o,a\}.\end{split}$
Then we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
Here $\mathcal{H}_{1}$ own Q functions satisfying optimal Bellman equation with data from $\mathcal{D}_{m}$, while $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ own Q functions satisfying optimal Bellman function with data from world models, which are trained with the single dataset and all datasets, respectively. Consequently, $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$ utilizes the generalization ability of world models, which extends the results in the fixed-task setting (Young et al. 2023). Moreover, $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$ illustrate similar dynamic structures across different tasks benefit learning general world models. Consequently, training general world models can narrow down the hypothesis of possible Q function and benefit agent training.
As existing world models (Hafner et al. 2019b, a) are mainly designed for the single-task setting, we propose reward-informed world models for the task-distribution setting. We first analyze the probabilistic graphical model of the task-distribution setting shown in Fig. 2, where rewards do not only rely on the previous states and actions but also the current task. Thus we can calculate the joint distribution as
$p(s_{1:T},o_{1:T},r_{1:T},a_{1:T-1},\mathcal{M})$ (1)
$=$ $p(\mathcal{M})\prod_{t=1}p(s_{t+1}|s_{t},a_{t})p(o_{t}|s_{t})p(r_{t}|s_{t},\mathcal{M}),$
and we choose the reward-informed inference model (the necessity of reward-informed policy in task generalization is discussed in Sec. 4.3) to approximate state posteriors as
$q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})=\prod_{t=1}^{T}q(s_{t}|s_{t-1},a_{t-1},r_{t-1},o_{t}).$ (2)
Based on this inference model, we can construct the variational lower bound of the log-likelihood on the data as
$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ \geq&\sum_{t=1}^{T}\mathbb{E}_{q(s_{t}|o_{\leq t},a_{<t},r_{<t})}\left[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})\right.\\ -&\left.\text{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)\right]\\ +&\mathbb{E}_{q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})}\left[\ln p(\mathcal{M}|s_{1:T})\right].\end{split}$ (3)
This result is a general form of the single-task setting (Hafner et al. 2019b) and its derivation is in Appendix A.2. The first two terms in Eq.(LABEL:eq_elbo) are for reconstructing observations, predicting rewards, and inferring states, which are similar to the single-task setting. The last novel term is dedicated to predicting the current task from historical information, which is beneficial for improving the generalization as it encourages inferring the current task context. Extending RSSM via Eq.(LABEL:eq_elbo), our reward-informed world models consist of:
Deterministic state model: $h_{t}=f(h_{t-1},s_{t-1},a_{t-1},r_{t-1}),$
Transition model: $p_{\theta}(s_{t}|h_{t}),$
Observation model: $p_{\theta}(o_{t}|h_{t},s_{t}),$
Reward model: $p_{\theta}(r_{t}|h_{t},s_{t}),$
Task model: $p_{\theta}(\mathcal{M}|h_{t},s_{t}).$
Here hidden state $h_{t}$ encodes historical states, actions, and rewards, by using gated recurrent unit (GRU) (Chung et al. 2014) as $f(\cdot)=\text{GRU}(\cdot)$. Then the transition model, observation model, reward model, and task model further predict state, observation, reward, and task context respectively.
4.2 Optimization
Based on the above analyses, we now introduce the training of TAD in detail. Following previous world models (Hafner et al. 2019a), we adopt an alternating training approach between the reward-informed world models and the policy.
To balance different tasks, when collecting data, TAD utilizes $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$ to store trajectories sampled from $\{\mathcal{M}_{m}\}_{m=1}^{M}$, respectively. Then, TAD samples data from each replay buffer and trains the reward-informed world models via the optimization objective following Eq. (LABEL:eq_elbo) as
$\begin{split}&L_{\text{TAD}}=\sum_{i=1}^{M}\mathbb{E}_{q}\left[\sum_{t=1}^{T}\ln p_{\theta}(o_{t}^{i}|h_{t}^{i},s_{t}^{i})+\sum_{t=1}^{T}\ln p_{\theta}(r_{t}^{i}|h_{t}^{i},s_{t}^{i})\right.\\ &-\left.\sum_{t=1}^{T}D_{\text{KL}}(q(s_{t}^{i}|h_{t}^{i},o_{t}^{i})\|p_{\theta}(s_{t}^{i}|h_{t}^{i}))+L_{\text{task}}\right].\end{split}$ (4)
In Eq. (LABEL:eq_loss), the first three items are similar to Dreamer for reconstructing observations, predicting rewards, and inferring states. Besides them, TAD includes an additional task term $L_{\text{task}}$ for predicting different tasks and learning task-aware embedding. In detail, we provide two alternatives: cross-entropy and self-contrastive, which are introduced below (We theoretically demonstrate their effectiveness in Sec. 4.3).
Cross-Entropy.
The last term in Eq. LABEL:eq_elbo indicates that we need to maximize the log probability of the task context to distinguish different tasks via historical information. Thus TAD-CE directly maximizes the log probability over different tasks represented by one-hot vectors following previous works (Yang et al. 2020) and $L_{\text{task}}$ is set as below.
$\begin{split}L_{\text{task}}=\sum_{t=1}^{T}\ln p_{\theta}(m^{i}|h_{t}^{i},s_{t}^{i}).\end{split}$ (5)
Supervised-Contrastive.
Besides directly maximizing the log probability, we further propose TAD-SC, borrowing the idea of supervised contractive learning (Khosla et al. 2020). In detail, TAD-SC keeps the task embeddings of the same task closer and the task embeddings of different tasks far apart, which benefits reward-informed world models to better distinguish different tasks. Formally, we assume that the task model maps all sampled data as $\{m_{j}\}_{j=1}^{M\times T}$ and set $L_{\text{task}}$ as
$\begin{split}L_{\text{task}}=\sum_{j=1}^{M\times T}\sum_{a\in A(j)}\ln\frac{\exp(m_{j}\cdot m_{a}/\tau)}{\sum_{b\neq j}\exp(m_{j}\cdot m_{b}/\tau)},\end{split}$ (6)
here $A(j)$ is the set of indices that are sampled with the same task of $m_{j}$, and $\tau$ is the temperature parameter, which is set as 0.1 following previous works (Khosla et al. 2020).
In terms of training the actor-critic that are parameterized neural networks, we extend the actor-critic learning in Dreamer to the task distribution. We first sample a series of states from the replay buffer and start from them to imagine trajectories via our reward-informed world models (results are in Fig. 3), which can capture invariant features and are beneficial for the agent to gain better generalization. After obtaining imagined trajectories, the actor-critic is optimized via maximizing the $\lambda$-return (Schulman et al. 2015) and regressing the TD targets (Sutton and Barto 2018), respectively.
Algorithm 1 Task Aware Dreamer (TAD)
0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize parameters of world models, the policy, and the critic.
[…]
2:     for update $\text{update step}=1,2,...,U$ do
3:        Sample $o$-$a$-$r$ pairs $\{(o_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}\}$ form each replay buffer $\mathcal{D}_{i},i=1,2,...,M$
4:        Calculate the deterministic state $h$ and further calculate model states $s$.
5:        Update the world models via optimizing Eq. (LABEL:eq_loss).
6:        Collect imagined trajectories from each $s$ via the policy and the world models and use these imagined trajectories to update the policy and the critic.
7:     end for
8:     Collect trajectories from $\mathcal{M}_{m}(m=1,2,...,M)$ and store them into the replay buffer $\mathcal{D}_{m}$.
9:  end for
10:  Evaluate the agent in testing environments $\{\mathcal{M}_{M+n}\}$.
4.3 Theoretical Analyses
Below, we provide theoretical analyses to show that TAD’s components are simple but effective for task generalization.
Are policies that utilize all historical information in TAD necessary for task generalization?
Below we introduce $3$ types of widely used policy hypotheses and show that $\Pi_{3}$, used in TAD, is necessary for handling task generalization.
  1. 1.
    Markovian policy set $\Pi_{1}$ (Sutton and Barto 2018; Yarats et al. 2021), i.e., $\Pi_{1}=\{\pi|\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})\}$, here $\Delta(\mathcal{A})$ represents a distribution over $\mathcal{A}$, which is widely used and optimal for the single-task setting;
  2. 2.
    $\mathcal{S}$-$\mathcal{A}$ memorized policy set $\Pi_{2}$ (Hafner et al. 2019a, 2020; Lee et al. 2020), i.e., $\Pi_{2}=\{\pi|\pi:\mathcal{H}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{H}=\cup_{t=1}^{\infty}\mathcal{H}_{t},\mathcal{H}_{t}=(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}$;
  3. 3.
    $\mathcal{S}$-$\mathcal{A}$$-\mathcal{R}$ memorized policy set $\Pi_{3}$ (Zintgraf et al. 2019; Rimon et al. 2024), i.e., $\Pi_{3}=\{\pi|\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A})\}$, here $\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$.
As illustrated in Fig. 1, naturally $\Pi_{1}\subseteq\Pi_{2}\subseteq\Pi_{3}$. Now we will analyze their expressive ability under task distribution $\mathcal{T}$. Denote $J_{\mathcal{T}}^{*}\triangleq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi}J_{\mathcal{M}}(\pi)\right]$ as the optimal return under $\mathcal{T}$, and $J_{\mathcal{T}}^{i}\triangleq\max_{\pi\in\Pi_{i}}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\right]$ as the optimal return for $\Pi_{i}(i=1,2,3)$ under $\mathcal{T}$. Our first result shows that, although $\Pi_{1}\subseteq\Pi_{2}$, they own the same expressive ability, i.e., $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}$, and are both sub-optimal:
Theorem 2 (Sub-Optimality of $\Pi_{1},\Pi_{2}$. Proof in Appendix A.3).
We set ¯$\bar{\mathcal{M}}=(\mathcal{S},\mathcal{A},\mathcal{P},\bar{\mathcal{R}},\gamma)$, here ¯$\bar{\mathcal{R}}=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}]$. For $\forall\pi\in\Pi_{2}$, we have $\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\bar{\mathcal{M}}}(\pi)$ and further $J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}\leq J_{\mathcal{T}}^{*}$.
Theorem 2 reveals that the cumulative returns of policies in $\Pi_{1}$ and $\Pi_{2}$ are the same as their returns in the “average” MDP $\bar{\mathcal{M}}$, where the reward function is the average of reward functions in different tasks. Also, as $\Pi_{1}$ and $\Pi_{2}$ only choose actions via current state or historical state-action pairs, they cannot distinguish different tasks and are both sub-optimal.
To quantitatively analyze the characteristic of $\mathcal{T}$ and the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$, we propose a novel metric Task Distribution Relevance (TDR) of the distribution $\mathcal{T}$ as
Definition 1 (Task Distribution Relevance).
For any task distribution $\mathcal{T}$ and state $s$, the Task Distribution Relevance of $\mathcal{T}$ and $s$ is defined as
$\begin{split}D_{\text{TDR}}(\mathcal{T},s)&=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\max_{a}Q_{\mathcal{M}}^{*}(s,a)]\\ &-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[Q_{\mathcal{M}}^{*}(s,a)\right].\end{split}$ (7)
Intuitively, TDR describes the relevance of $\mathcal{T}$ via optimal $Q$ functions, which determine the distribution of optimal actions in corresponding tasks. Based on TDR, we can bound the gap:
Theorem 3 (Proof in Appendix A.4).
Assume $\pi_{\mathcal{M}}^{*}=\mathop{\arg\max}_{\pi}J_{\mathcal{M}}(\pi)$, for $\forall\pi\in\Pi_{1}$, we have
$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}[D_{\text{TDR}}(\mathcal{T},s)]\end{split}$ (8)
Thus $J_{\mathcal{T}}^{*}-J_{\mathcal{T}}^{2}=J_{\mathcal{T}}^{*}-J_{\mathcal{T}}^{1}\geq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right]$, here $\pi^{*}=\arg\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)$.
Theorem 3 demonstrates that the gap between $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ and $J_{\mathcal{T}}^{*}$ is related to the TDR of $\mathcal{T}$. When considering $\mathcal{T}$ with high TDR, i.e., the optimal $Q$ values in different tasks differ greatly, the performance of $\Pi_{1}$ and $\Pi_{2}$ will be extremely poor since they cannot differentiate different tasks and their expressive abilities are significantly limited. This conclusion is further verified empirically in the experimental section. Moreover, we can show that $J_{\mathcal{T}}^{1},J_{\mathcal{T}}^{2}$ might be arbitrarily small when $J_{\mathcal{T}}^{3}$ might be arbitrarily close to $J_{\mathcal{T}}^{*}$ (More details are in Appendix A.5), which demonstrates that $\Pi_{3}$ owns stronger expressive ability. Consequently, it is necessary to utilize $\Pi_{3}$ to distinguish different tasks for enhancing expression ability and generalization over the task distribution.
Is optimizing $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ helpful for task generalization?
Now we will show that, although $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ is a simple term, optimizing it can be effective for reaching the generalizable agent in the task distribution.
As the input space of $\Pi_{3}$ is $\mathcal{L}$, i.e., all partial trajectories, our major result analyzes the relation between the policy $\pi\in\Pi_{3}$ and the task posterior $p(\mathcal{M}|l),l\in\mathcal{L}$. As $h_{t}$ encodes all historical information, $p_{\theta}(\mathcal{M}|h_{t},s_{t})$ can be rewritten as $p(\mathcal{M}|l),l\in\mathcal{L}$.
Theorem 4 (Informally, detailed analyses and proof are in Appendix A.6).
For any policy $\pi\in\Pi_{3}$, we have
$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl,\end{split}$ (9)
here $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
Consequently, maximizing $p(\mathcal{M}|l)$, i.e., making the distribution of $p(\mathcal{M}|l)$ to be closer to some Dirac distribution, can significantly reduce the right part of Eq. (LABEL:eq_pi3), thus is effective for improving the generalization ability of $\pi$. More details and discussion are also provided in Appendix A.6.
Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run__\__back
&balance__\__sparse &prostrate&flip &flip&flip__\__back
CURL $\Pi_{1}$ 994.5 $\pm$ 3.6 254.1 $\pm$ 9.2 229.7 $\pm$ 10.9
SAC+AE $\Pi_{1}$ 992.5 $\pm$ 2.6 256.9 $\pm$ 5.9 225.8 $\pm$ 10.1
PlaNet $\Pi_{2}$ 309.5 $\pm$ 59.9 606.7 $\pm$ 152.9 244.8 $\pm$ 17.8
Dreamer $\Pi_{2}$ 974.2 $\pm$ 5.8 722.2 $\pm$ 12.6 241.1 $\pm$ 19.5
MAMBA $\Pi_{3}$ 994.7 $\pm$ 3.2 436.8 $\pm$ 116.1 375.6 $\pm$ 44.0
TAD-CE (Ours) $\Pi_{3}$ 998.9 $\pm$ 0.4 778.9 $\pm$ 63.1 549.6 $\pm$ 28.6
TAD-SC (Ours) $\Pi_{3}$ 982.6 $\pm$ 2.0 807.8 $\pm$ 85.2 588.8 $\pm$ 20.2
[…]
Algorithms Hypothesis Cheetah__\__speed Pendulum__\__angle Walker__\__speed
Train Test Train Test Train Test
CURL $\Pi_{1}$ 211.7 $\pm$ 13.7 57.4 $\pm$ 26.6 140.2 $\pm$ 1.7 46.1 $\pm$ 29.8 127.0 $\pm$ 33.7 77.5 $\pm$ 11.5
SAC+AE $\Pi_{1}$ 182.2 $\pm$ 7.6 115.2 $\pm$ 10.1 130.6 $\pm$ 12.2 89.0 $\pm$ 25.7 136.8 $\pm$ 34.4 27.5 $\pm$ 10.5
PlaNet $\Pi_{2}$ 176.6 $\pm$ 25.9 83.0 $\pm$ 52.2 92.5 $\pm$ 31.3 70.6 $\pm$ 18.4 173.9 $\pm$ 19.3 58.4 $\pm$ 23.7
Dreamer $\Pi_{2}$ 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 87.8 $\pm$ 16.1 87.3 $\pm$ 20.5 197.6 $\pm$ 24.6 10.0 $\pm$ 6.5
MAMBA $\Pi_{3}$ 568.0 $\pm$ 229.1 475.7 $\pm$ 316.6 153.8 $\pm$ 34.7 121.1 $\pm$ 29.2 104.5 $\pm$ 25.7 99.9 $\pm$ 43.1
TAD-CE (Ours) $\Pi_{3}$ 937.4 $\pm$ 9.8 909.8 $\pm$ 21.9 283.9 $\pm$ 16.2 163.8 $\pm$ 53.0 241.2 $\pm$ 36.8 156.5 $\pm$ 129.6
TAD-SC (Ours) $\Pi_{3}$ 919.3 $\pm$ 21.9 906.7 $\pm$ 21.7 204.4 $\pm$ 62.2 143.9 $\pm$ 70.7 159.0 $\pm$ 36.9 104.3 $\pm$ 43.7
[…]
Algorithms Hypothesis Half-Cheetah-Fwd-Back(1e7) Half-Cheetah-Vel(1e7) Humanoid-Direc-2D(1e6)
Train&Test Train Test Train Test
PlaNet $\Pi_{2}$ 30.5 $\pm$ 42.9 -198.1 $\pm$ 1.9 -202.1 $\pm$ 1.8 215.9 $\pm$ 72.3 220.6 $\pm$ 75.3
Dreamer $\Pi_{2}$ 127.4 $\pm$ 181.8 -151.4 $\pm$ 0.4 -169.4 $\pm$ 1.2 260.5 $\pm$ 48.9 263.5 $\pm$ 52.3
RL2(zero-shot) $\Pi_{3}$ 1070.7 $\pm$ 109.7 -70.3 $\pm$ 6.7 191.9 $\pm$ 50.8
RL2(few-shot) $\Pi_{3}$ 1006.9 $\pm$ 26.4 -146.9 $\pm$ 0.4 268.8 $\pm$ 30.2
MAML(few-shot) 429.3 $\pm$ 81.4 -121.0 $\pm$ 37.1 205.3 $\pm$ 34.7
VariBAD(zero-shot) $\Pi_{3}$ 1177.5 $\pm$ 94.9 -58.4 $\pm$ 20.6 260.3 $\pm$ 61.6
TAD-CE (Ours) $\Pi_{3}$ 1455.8 $\pm$ 78.3 -49.3 $\pm$ 1.9 -47.1 $\pm$ 0.3 339.5 $\pm$ 78.7 335.5 $\pm$ 70.5
TAD-SC (Ours) $\Pi_{3}$ 1541.5 $\pm$ 114.8 -50.5 $\pm$ 1.6 -49.6 $\pm$ 1.6 260.2 $\pm$ 185.0 249.0 $\pm$ 168.9
[…]
5 Experiments
We now present empirical results to answer the following questions:
  • Can we verify the analyses about TDR, i.e., the expressive abilities of $\Pi_{1},\Pi_{2}$ are severely restricted in task distributions with high TDR? (Sec. 5.2)
  • How about TAD’s generalization ability when handling image-based and state-based observations? (Sec. 5.3)
  • Can TAD be extended to more general settings like dynamic generalization? (Sec. 5.4)
5.1 Experimental Setup
Image-based Control.
To verify our analyses of TDR in Sec. 4.3, we consider several task combinations in DeepMind Control suite (DMC) (Tassa et al. 2018): (1) Cartpole-balance&balance__\__sparse, which shares the same optimal actions with TDR 0; (2) two task combinations with non-zero TDR of Walker-stand&walk&prostrat&flip, and Cheetah-run&run__\__back&flip&flip__\__flip__\__back, which are widely used in multi-task unsupervised RL (Sekar et al. 2020; Laskin et al. 2021). For example, Cheetah-run and Cheetah-flip hope a two-leg robot to move forward by running and flip around the torso, respectively (see Fig. 3), yielding almost opposite optimal Q functions with huge TDR. In Appendix C.1, we introduce more details about these task combinations.
To evaluate generalization with image-based observations, we extend tasks in DMC and design three task distributions: (1) Cheetah__\__speed($\alpha,\ \beta$), which extends Cheetah-run and hopes the agent to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$; (2) Pendulum__\__angle($\alpha,\ \beta$), extending Pendulum-swingup to keep pendulum’s pole within the target angle interval $(\arccos\alpha,\ \arccos\beta)$; and (3) Walker__\__speed($\alpha,\ \beta$), which is based on Walker-run and requires the planar walker to run within the target speed interval $(\alpha-\beta,\ \alpha+\beta)$. For each task distribution, we sample $4$ training tasks and $2$ additional test tasks. More details are in Appendix C.2.
State-based Control.
To demonstrate the scalability of TAD, we also consider some state-based continuous robotic control task distributions simulated via MuJoCo (Todorov, Erez, and Tassa 2012). Following previous work (Finn, Abbeel, and Levine 2017), we choose tasks distributions: (1) Half-Cheetah-Fwd-Back, which owns two opposite tasks; and (2) Half-Cheetah-Vel and Humanoid-Direc-2D, which are task distributions with 100 training tasks 30 test tasks. More details of these task distributions are in Appendix D.
Baselines.
In DMC, we choose two model-free methods employing Markovian policies $\Pi_{1}$: CURL (Laskin, Srinivas, and Abbeel 2020) and SAC+AE (Yarats et al. 2021). In addition, we choose two classic world models, PlaNet (Hafner et al. 2019b) and Dreamer (Hafner et al. 2019a), which utilize historical state-actions in policies and belong to $\Pi_{2}$. Moreover, we take a SOTA model-based meta RL method belonging to $\Pi_{3}$: MAMBA (Rimon et al. 2024). In state-based control, besides PlaNet and Dreamer, we take some meta RL methods like MAML (Finn, Abbeel, and Levine 2017), RL2 (Duan et al. 2016), and VariBAD (Zintgraf et al. 2019), including zero-shot and few-shot evaluation, as reference.
Metrics.
For the task combinations, we evaluate the average return of all tasks to verify TDR. For the task generalization settings, we train agents in training tasks and evaluate their generalization abilities in test tasks. For all experiments, we repeat 5 different random seeds and report the mean $\pm$ std to mitigate the effects of randomness following previous works (Hafner et al. 2019a; Rimon et al. 2024).
Figure 3: Sampled trajectories and imaginary trajectories of TAD for different tasks (Cheetah-run and Cheetah-flip).
Figure 4: The t-SNE clustering of state embeddings for different tasks sampled via Dreamer, TAD-CE, and TAD-SC.
5.2 Experimental Results for TDR
To validate our analyses of TDR, we report different algorithms’ performance of different task combinations in Table 1. Results show that TAD-CE and TAD-SC outperform all baselines, especially in those environments with high TDR. As derived in Theorem 3, in task combinations where TDR is 0 like Cartpole-balance&balance__\__sparse, different tasks share the same optimal action and $\Pi_{1},\Pi_{2}$ own the optimal policy, thus baselines like CURL and Dreamer perform well. Conversely, in other task combinations, different tasks’ optimal Q functions may differ a lot (like Chetah-run and Cheetah-run__\__back) and TDR is significantly huge. Thus methods with policies in $\Pi_{1},\Pi_{2}$ like CURL and Dreamer can not differentiate different tasks and perform poorly. Aligning with Theorem 3, baselines utilizing $\Pi_{3}$ like MAMBA outperform other baselines and TAD improves performance conspicuously.
Moreover, we present some visualization to better understand how TAD works. Fig, 3 shows video predictions of TAD for different tasks (Cheetah-run and Cheetah-flip). We use the same agent trained by TAD to sample trajectories for these two tasks and show their trajectories in lines 1 and 3 respectively. Given the first 7 steps as the context (their observations are similar but the rewards are different), we directly imagine the future 55 steps by our trained reward-informed world model, of which the results are in lines 2 and 4 of Fig. 3 respectively. As shown here, TAD performs differently to handle different tasks only by receiving different rewards. Moreover, our reward-informed world models demonstrate long-term prediction and high-quality reconstruction capabilities, which reveals the potential for training future large-scale world models. In Fig. 4, we sample trajectories from agents trained by Dreamer, TAD-CE, as well as TAD-SC, and visualize the states of different tasks, of which the dimensions are reduced for visualization by t-SNE (Van der Maaten and Hinton 2008). As shown here, states in TAD-CE and TAD-SC of different tasks are clearly distinguished, while Dreamer can not differentiate them and perform the same in different tasks. This result demonstrates TAD can effectively learn task-aware information and distinguish different tasks. Videos of different tasks are provided in supplementary materials.
5.3 Experimental Results for Task Generalization
To answer the second question, we report the task generalization results of image-based and state-based environments in Table 2-3. In Table 2, TAD shows more powerful generalization abilities than all baselines. This demonstrates that TAD can both handle multiple training tasks simultaneously and generalize to unseen test tasks effectively. We also provide some visualization results in Appendix C.3 with videos in supplementary materials.
Additionally, in Table 3, we compare TAD with existing meta RL methods in state-based environments, of which the training timesteps are 1e7, 1e7, and 1e6, respectively. Many context-based meta RL methods, including RL2, PEARL, and VariBAD, utilize historical rewards and belong to $\Pi_{3}$, thus they can distinguish different tasks, which also verify our analyses. As shown in Table 3, TAD achieves a significant improvement for both training tasks and test tasks, since TAD is aware of task information for generalizing to unseen tasks.
5.4 Ablation Study
Reward Signals.
In Fig. 5, we do ablation studies about the task term in TAD, i.e., we consider Dreamer(w/ r) that only integrates $\Pi_{3}$ into Dreamer. Results show that just providing rewards can help Dreamer distinguish different tasks, which verifies our analyses in TDR. Also, TAD-CE and TAD-SC show superior performance, demonstrating that our task model and corresponding ELBO are effective for task generalization. Also, we do ablation studies on reward signals for model-free methods like CURL in Appendix E.1.
Figure 5: Ablation study on Reward Signal.
Extension to Dynamic Generalization.
To answer the third question, we evaluate TAD in more general settings with different observations, dynamics, and/or actions. As TAD utilizes all historical information to infer the environment, it can be directly applied to these settings. We design task distribution with different embodiments (Acrobot-Cartpole-Pendulum, Walker-Cheetah-Hopper), as well as different dynamics (Cheetah-run__\__mass, Walker-walk__\__mass). More details about environments and results are in Appendix E.2, where TAD achieves much greater performance compared to baselines and show potential in further handling dynamic generalization and even cross-embodiment tasks.
5.5 Limitations and Discussion
In terms of limitations, TAD assumes that the task context is continuously related to historical information for generalizing to unseen tasks. Thus TAD might be difficult to generalize in sparse-reward settings. We further demonstrate that without extra knowledge or finetuning, zero-shot generalization to unseen tasks with extremely sparse rewards is impossible since there is no way to distinguish different tasks (Appendix A.7). Fortunately, in relatively sparse reward settings, we conduct experiments in Appendix E.3 to show that TAD can infer the current task and generalize to unseen tasks well.
6 Conclusion
In this work, we propose a novel framework of TAD that handles different tasks via all historical information and utilizes novel reward-informed world models to capture invariant latent features. In TAD, we calculate the corresponding variational lower bound of the data log-likelihood, which includes a novel loss term to distinguish different tasks via states. To explain components in TAD, we introduce a novel metric TDR to capture the relevance of the task distribution and show that Markovian policies perform poorly in tasks with high TDR. Experiments in image-based and state-based settings demonstrate that TAD can remarkably improve the performance of handling different tasks meanwhile, especially for high TDR ones, and successfully generalize to unseen tasks.
Appendix A Proof of Theorems
In this section, we will provide detailed proofs of theorems.
A.1 The Proof of Theorem 1
Proof.
First, we will show that $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$, following the proof in (Young et al. 2023) that considers the setting that the rewards of all environments are the same.
$\begin{split}&\{q_{m}\}\in\mathcal{H}_{2}\\ \Leftrightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\\ \Rightarrow&\exists(r_{m})_{m=1}^{M}::p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o_{t}^{m},a_{t}^{m}),a^{\prime}),\forall m,t\\ \Rightarrow&q_{m}(o_{t}^{m},a_{t}^{m})=r^{m}_{t}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime}),\forall m,t\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{1},\end{split}$ (10)
thus we have $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$. Next, we prove that $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$, which is mainly because we can utilize similar dynamic structures from different tasks to narrow down the hypothesis spaces of the dynamic model.
$\begin{split}&\{q_{m}\}\in\mathcal{H}_{3}\\ \Leftrightarrow&\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall m,o,a\\ \Rightarrow&\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\forall m,t\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall m,o,a\\ \Leftrightarrow&\{q_{m}\}\in\mathcal{H}_{2},\end{split}$ (11)
thus we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$. ∎
A.2 The Derivation of the ELBO
We use ${\color[rgb]{0,139,69}q}$ to represent $q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})$, ${\color[rgb]{0,0,1}\hat{q}}$ to present $q(s_{t}|o_{\leq t},a_{<t},r_{<t})$, ${\color[rgb]{1,0,0}\tilde{q}}$ to represent $q(s_{t-1}|o_{\leq(t-1)},r_{<(t-1)},a_{<(t-1)})$, and we have
$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T},\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(o_{1:T},r_{1:T}|\mathcal{M},s_{1:T})p(\mathcal{M}|s_{1:T})\right]\\ =&\ln\mathbb{E}_{p(s_{1:T}|a_{1:T})}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\right]\\ =&\ln\mathbb{E}_{{\color[rgb]{0,139,69}q}}\left[p(\mathcal{M}|s_{1:T})\prod_{t=1}^{T}p(o_{t},r_{t}|\mathcal{M},s_{t})\frac{p(s_{t}|s_{t-1},a_{t-1})}{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}\right]\\ \geq&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})+\sum_{t=1}^{T}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ +&\mathbb{E}_{{\color[rgb]{0,139,69}q}}\sum_{t=1}^{T}\left[\ln p(s_{t}|s_{t-1},a_{t-1})-\ln q(s_{t}|o_{\leq t},r_{<t},a_{<t})\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]\\ +&\sum_{t=1}^{T}\left[\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln p(o_{t},r_{t}|\mathcal{M},s_{t})-\mathbb{E}_{{\color[rgb]{0,139,69}q}}\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}{\color[rgb]{1,0,0}\tilde{q}}}\left[\ln\frac{q(s_{t}|o_{\leq t},r_{<t},a_{<t})}{p(s_{t}|s_{t-1},a_{t-1})}\right]\\ =&\mathbb{E}_{{\color[rgb]{0,139,69}q}}[\ln p(\mathcal{M}|s_{1:T})]+\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{0,0,1}\hat{q}}}[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})]\\ -&\sum_{t=1}^{T}\mathbb{E}_{{\color[rgb]{1,0,0}\tilde{q}}}[\mathrm{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)].\end{split}$ (12)
Thus we have proven the ELBO.
A.3 The Proof of Theorem 2
Proof.
We first prove that for $\forall\pi\in\Pi_{2}$, we have $\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\bar{\mathcal{M}}}(\pi)$.
For any $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\gamma)\sim\mathcal{T}$, we can use the policy $\pi$ to interact with $\mathcal{M}$ and get the trajectory $\tau=(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},r_{1}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},r_{2}^{\mathcal{M}},...)$. Since the dynamic transition $\mathcal{P}$ is the same for all $\mathcal{M}$ and the policy $\pi\in\Pi_{2}$ only depends on historical states and actions, we naturally have that the distribution of all states and actions $(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},...)$ are the same for all $\mathcal{M}\sim\mathcal{T}$ as well as $\bar{\mathcal{M}}$. Consequently, we have
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]\\ =&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\mathcal{M}}(\tau)]=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{\mathcal{M}}\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[r_{t}^{\mathcal{M}}]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}[\bar{\mathcal{R}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\bar{\mathcal{M}}}(\tau)]=J_{\bar{\mathcal{M}}}(\pi).\end{split}$ (13)
It is well known that the optimal policy in single MDP is memory-less, i.e., $\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)$. Consequently, we have
$\begin{split}J_{\mathcal{T}}^{2}&=\max_{\pi\in\Pi_{2}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)\\ &=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\mathcal{T}}^{1}\\ &\leq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)\right]=J_{\mathcal{T}}^{*}.\end{split}$ (14)
Thus we have proven this result. ∎
A.4 The Proof of Theorem 3
Proof.
Our proof follows some previous work (Kakade and Langford 2002; Ying et al. 2022). First, we consider the bellman equation of value function of $\pi,\pi_{\mathcal{M}}^{*}\in\Pi_{1}$ in $\mathcal{M}$ as
$\begin{split}V_{\mathcal{M},\pi}(s)=&\sum_{a}\pi(a|s)\left[\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi}(s^{\prime})\right],\\ V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)=&\sum_{a}\pi_{\mathcal{M}}^{*}(a|s)\left[\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\right].\end{split}$
Defining $\Delta V(s)\triangleq V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)$ as the difference of these two value functions, we can further deduce that
$\begin{split}&V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)\\ =&\gamma\sum_{a}\Delta\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})+\sum_{a}\Delta\pi(a|s)\mathcal{R}(s,a)\\ =&\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime}),\end{split}$ (15)
here $\Delta\pi(a|s)=\pi(a|s)-\pi_{\mathcal{M}}^{*}(a|s)$. Since Eq. (LABEL:15) holds for any $s$, thus we calculate its expectation for $s\sim d_{\mathcal{M}}^{\pi_{\mathcal{M}}^{*}}$:
$\begin{split}&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\Delta V(s)\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)[V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)]\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\mathcal{P}(s^{\prime}|s,a)\right].\end{split}$ (16)
Since $\gamma\sum_{s^{\prime}}d_{\mathcal{M}}^{\pi}(s^{\prime})\sum_{a}\pi(a|s^{\prime})\mathcal{P}(s|s^{\prime},a)=d_{\mathcal{M}}^{\pi}(s)-(1-\gamma)\mathcal{P}(s_{0}=s)$, we have
$\begin{split}\sum_{s}d_{\mathcal{M}}^{\pi}(s)&\Delta V(s)=\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[d_{\mathcal{M}}^{\pi}(s^{\prime})-(1-\gamma)\mathcal{P}(s_{0}=s^{\prime})\right].\end{split}$ (17)
By moving the second term of the right part in Eq. (17) to the left part, we can deduce that
$\begin{split}&(1-\gamma)\sum_{s^{\prime}}\Delta V(s^{\prime})\mathcal{P}(s_{0}=s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a),\end{split}$ (18)
thus we can calculate that
[…]
Consequently, we have
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\int_{a}\pi(a|s)Q_{\mathcal{M}}^{*}(s,a)da\right]\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}Q_{\mathcal{M}}^{*}(s,a)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}\left[D_{\text{TDR}}(\mathcal{T},s)\right].\end{split}$ (20)
Since $J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})=\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)$, we have
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)-\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right],\end{split}$ (21)
Thus we have proven this result. ∎
A.5 Expressive Ability of $\Pi_{3}$
Proposition 1.
For $\forall\epsilon_{1},\epsilon_{2}$ satisfying $0<\epsilon_{1}\leq 1,0<\epsilon_{2}\leq 1$, there exists a task distribution $\mathcal{T}$ satisfying that
$\begin{split}J_{\mathcal{T}}^{1}=J_{\mathcal{T}}^{2}\leq\epsilon_{1},~{}~{}J_{\mathcal{T}}^{3}\geq 1-\epsilon_{2},~{}~{}J_{\mathcal{T}}^{*}=1.\end{split}$ (22)
Proof.
Given fixed discount factor $\gamma\in(0,1)$, we first take $n\in\mathbb{N}$ satisfying $n\geq\frac{1}{\epsilon_{1}}$, $A=\epsilon_{2}\frac{n}{n-1}\frac{1-\gamma}{1-\gamma^{n}}$, and $B=\frac{1-\gamma}{\gamma^{n+1}}\left(1-\epsilon_{2}\frac{n}{n-1}\right)$. We construct state sets $\mathcal{S}=\{s_{t}^{l}\}(t=0,1,...,\infty,l=1,...,n)$ and action sets $\mathcal{A}=\{a_{j}\}_{j=1}^{n}$. Then we construct $n$ tasks $\mathcal{M}_{i}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{i},\gamma),i=1,...,n$, which share the same dynamic $\mathcal{P}$ and different reward functions $\mathcal{R}_{i}$. The initial state of each task is $s_{0}^{1}$, and the dynamic as well as reward functions are as below
$\begin{split}&\mathcal{P}(s_{t}^{l},a_{j})=\mathbb{I}(s=s_{t+1}^{j}),\quad j,l=1,...,n;\ t=0,1,...,\infty\\ &\mathcal{R}_{i}(s_{t}^{l},a_{j})=f(t)\mathbb{I}(i=j),\quad i,j,l=1,...,n;\ t=0,1,...\infty\\ &f(t)=\left\{\begin{array}[]{ll}A,&t\leq n-1\\ B,&t\geq n\\ \end{array}\right.\end{split}$ (23)
Take $\mathcal{T}$ as the uniform distribution over $\mathcal{M}_{1},...,\mathcal{M}_{n}$, thus we have
$\begin{split}J_{\mathcal{M}_{i}}^{*}=&A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...\\ =&A\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1\\ J_{\mathcal{T}}^{*}=&\frac{1}{n}\sum_{i=1}^{n}J_{\mathcal{M}_{i}}^{*}=1.\\ \end{split}$ (24)
Since our construction satisfies $\mathbb{E}_{\mathcal{T}}[\mathcal{R}_{i}(s_{k}^{l},a_{j})]=\frac{f(k)}{n}$, for $\forall\pi\in\Pi_{2}$, we have
$\begin{split}J_{\mathcal{T}}(\pi)=&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...)\\ =&\frac{1}{n},\\ J_{\mathcal{T}}^{1}=&J_{\mathcal{T}}^{2}=\max_{\pi\in\Pi_{2}}J_{\mathcal{T}}(\pi)=\frac{1}{n}\leq\epsilon_{1}.\end{split}$ (25)
Moreover, we can construction an agent $\hat{\pi}\in\Pi_{3}$ that takes action via the historical trajectory $\hat{\tau}_{t}=(\hat{s}_{0},\hat{a}_{0},\hat{r}_{0},...\hat{s}_{t})$:
$\begin{split}\hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=t+1),\quad t=0,1,...,n-1\\ \hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=i),\quad t=n,...,\infty\\ \end{split}$ (26)
here $i=\arg\max\{\hat{r}_{0},\hat{r}_{1},...,\hat{r}_{n-1}\}+1$, thus we have
$\begin{split}&J_{\mathcal{T}}^{3}\geq J_{\mathcal{T}}(\hat{\pi})\\ =&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1})+B\gamma^{n}+B\gamma^{n+1}+...\\ =&\frac{A}{n}\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1-\frac{(n-1)A}{n}\frac{1-\gamma^{n}}{1-\gamma}\\ \geq&1-\epsilon_{2}.\end{split}$ (27)
Thus we have proven this result. ∎
A.6 Proof and Discussion of Theorem 4
In this part, we introduce an informed version of Theorem 4, about why optimizing $p(\mathcal{M}|l)\forall l\in\mathcal{L}$ is beneficial for task generalization, with detailed proofs. Recall that we consider the policy hypothesis $\mathcal{H}_{3}$ here, that each policy $\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A}),\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$. As directly such $\mathcal{S}-\mathcal{A}-\mathcal{R}$ memorized policy is difficult, we consider an alternative MDP as ~$\tilde{\mathcal{M}}=(\mathcal{L},\mathcal{A},\mathcal{P}_{\mathcal{M}},\mathcal{R}_{\mathcal{M}},\gamma)$. For $\forall l=(s_{1},a_{1},r_{1},...,s_{t})\in\mathcal{L}_{t}\subseteq\mathcal{L}$, we can sample the action $a_{t}$ from the distribution $\pi(\cdot|l)$. Then the environment will feedback the reward signal $r_{t}=\mathcal{R}_{\mathcal{M}}(l,a_{t})=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$, we can sample $s_{t+1}$ from the distribution $\mathcal{P}(\cdot|s_{t},a_{t})$, and the next environment state will be $l^{\prime}=(s_{1},a_{1},r_{1},...,s_{t},a_{t},r_{t},s_{t+1})\in\mathcal{L}_{t+1}$. In summary, we set the distribution $p(l^{\prime}|l,a_{t})$ as the corresponding dynamic $\mathcal{P}_{\mathcal{M}}$. Notice that although all $\mathcal{M}\in\mathcal{T}$ shares the same dynamic $\mathcal{P}$, their new dynamic $\mathcal{P}_{\mathcal{M}}$ are different since the dynamic is related to the given reward. An obvious advantage of introducing $\tilde{\mathcal{M}}$ is that $\pi\in\Pi_{3}$ is now Markovian in $\tilde{\mathcal{M}}$ and it is much easier to analyze its performance.
Obviously, we can set $J_{\mathcal{M}}(\pi)=J_{\tilde{\mathcal{M}}}(\pi),Q_{\mathcal{M}}^{*}(s,a)=Q_{\tilde{\mathcal{M}}}^{*}(s,a)$ to simplify the notation, and we can prove that
Theorem 5.
For any policy $\pi\in\Pi_{3}$, we have
$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (28)
here $p(l)$ is a distribution of $\mathcal{L}$ related to $\mathcal{T},\pi$ and $p(\mathcal{M}|l)$ is the task posterior related to $\pi$.
Proof.
$\forall\pi\in\mathcal{H}_{3}$, as $\pi$ is Markovian in $\tilde{\mathcal{M}}$, we can directly utilize the proof of Theorem 3 from the beginning to Eq. (LABEL:app_eq_11), and the only difference is that the dynamics in $\tilde{\mathcal{M}}$ are different but the dynamics in $\mathcal{M}$ are the same. Thus we need to change the Eq. (LABEL:app_eq_12) as
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{l\sim d_{\tilde{\mathcal{M}},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int p(\mathcal{M})\int_{\mathcal{L}}d_{\tilde{\mathcal{M}},\pi}(l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\int p(\mathcal{M}|l)\left[\max_{a}Q_{\mathcal{M}}^{*}(l,a)\right.\\ &\left.-\int_{a}\pi(a|l)Q_{\mathcal{M}}^{*}(l,a)da\right]\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (29)
[…]
Finally, we will show that maximizing $p(\mathcal{M}|l)$ is helpful for task generalization. In the training stage, we will sample a task $\mathcal{M}$ and corresponding state-action-reward pairs $l$, thus optimizing $p(\mathcal{M}|l)$ will make it to be closer to some Dirac distribution. In such a situation, for each $l$, we can infer a “most possible” posterior task $\mathcal{M}_{l}$ with high $p(\mathcal{M}_{l}|l)$, thus we can approximately take $\pi(l)=\arg\max_{a}Q_{\mathcal{M}_{l}}^{*}(l,a)$ and the optimal gap calculated by Eq. (LABEL:app_ep_20) will be controlled.
A.7 Sparse Reward Task
In this part, we show that generalizing to tasks with the same dynamics and sparse rewards without extra knowledge (like context) is extremely difficult and sometimes impossible. It is because we cannot distinguish them via historical information. Here we construct an example.
Assume there are $n$ MDPs, $\mathcal{M}_{i}(i=1,...,n)$, each MDP share the same state set $\mathcal{S}=\{s_{1},...,s_{T}\}$ and action set $\mathcal{A}=\{a_{1},...,a_{n}\}$. The initial state is $s_{1}$ and the dynamic is that $\mathcal{P}(s_{t+1}|s_{t},a_{i})=1(t=1,...,T-1,i=1,...,n),\mathcal{P}(s_{T}|s_{T},a_{i})=1(i=1,...,n)$. As for the reward function, we define that
$\mathcal{R}_{\mathcal{M}_{i}}(s_{T-1},a_{i})=1,i=1,...,n.$ (30)
and the reward function is 0 otherwise. In this case, any policies (including markovian, state-action memorized, and state-action-reward memorized) in this task distribution can only handle one task since they can not distinguish them.
Appendix B Pseudo Code of TAD
The detailed pseudo code of TAD is provided in Algorithm 2.
Algorithm 2 Task Aware Dreamer (TAD)
0:  $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$, $M$ replay buffers $\{\mathcal{D}_{m}\}_{m=1}^{M}$, $N$ test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$, initialize neural network parameters of world models, the policy, and the critic
1:  while not converge do
2:     //$Model\ Training$
3:     for update $\text{update step}=1,2,...,U$ do
4:        Sample observation-action-reward pairs form each replay buffer $\{(o_{t}^{i},a_{t}^{i},r_{t}^{i})_{t=1}^{T}\}\sim\mathcal{D}_{i},i=1,2,...,M$
5:        Calculate the deterministic state $h$ and further calculate model states $s$.
6:        Update the world models via optimizing Eq. (6).
7:        Collect imagined trajectories from each $s_{t}$ via the policy and the world models.
8:        Use these imagined trajectories to update the policy and the critic.
9:     end for
10:     //$Data\ Collection$
11:     for $m=1,2,...,M$ do
12:        $o_{1}\leftarrow\mathcal{M}_{m}.reset()$
13:        for sample $\text{sample step}=1,2,...,S$ do
14:           Compute $h_{t},s_{t}$ and sample action $a_{t}$ via the policy.
15:           $r_{t},o_{t+1}\leftarrow\mathcal{M}_{m}.step(a_{t})$
16:        end for
17:        Add these data to the replay buffer $\mathcal{D}_{m}$.
18:     end for
19:  end while
20:  //$Model\ Evaluation$
21:  for $n=1,2,...,N$ do
22:     $o_{1}\leftarrow\mathcal{M}_{M+n}.reset()$
23:     while the environment not done do
24:        Compute $h_{t},s_{t}$ and sample action $a_{t}$ via the policy.
25:        $r_{t},o_{t+1}\leftarrow\mathcal{M}_{M+n}.step(a_{t})$
26:     end while
27:  end for
Appendix C Experimental Details for DMControl
C.1 Details of All Task Combinations
[…]
  • Cartpole-balance&sparse. This task combination includes two tasks: Cartpole-balance and Cartpole-balance__\__sparse, which both hope to balance an unactuated pole with dense and sparse rewards respectively. The optimal actions of these tasks are both hoped to balance the agent and thus TDR here is 0.
  • Walker-stand&walk&prostrate&flip. This task combination includes four tasks: Walker-stand, Walker-walk, Walker-prostrate, and Walker-flip. Walker-stand hopes the height of a two-leg robot to be larger than the target height. Walker-walk hopes the height of the improved planar walker to be larger than a target height and the speed of the robot to be larger than another target speed. Walker-prostrate hopes the height of the robot to be lower than the target height. Finally, Walker-flip hopes the robot to stand and move forward to the target speed by executing a rapid twist and jump. In detail, their reward functions can be roughly described as
    $\begin{split}\mathcal{R}_{\text{stand}}=&\text{tolerance}(\text{height},(1.2,\infty)),\\ \mathcal{R}_{\text{walk}}=&\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{speed},(1,\infty)).\\ \mathcal{R}_{\text{prostrate}}=&\text{tolerance}(\text{height},(0.0,0.2)).\\ \mathcal{R}_{\text{flip}}=&\text{tolerance}(\text{height},(1.2,\infty))\\ &*\text{tolerance}(\text{angmomentum},5,\infty).\end{split}$ (31)
    In this situation, for all states, the optimal actions of Walker-walk/stand and Walker-prostrate are opposite, and TDR here is huge.
  • Cheetah-run&run__\__back&flip&flip__\__back. This task combination includes four tasks: Cheetah-run, Cheetah-run__\__back, Cheetah-flip, and Cheetah-flip__\__back. Cheetah-run hopes to control a running planar biped to run forward within a target speed. Cheetah-run__\__back, differently, hopes to control the Cheetah robot to run backward within a target speed. Cheetah-flip hopes the robot to move forward to the target speed by executing a rapid twist and jump. Similarly, Cheetah-flip__\__back hopes to control the robot to move backward by flipping.
    $\begin{split}\mathcal{R}_{\text{Cheetah}\_\text{run}}=&\text{tolerance}(\text{speed},(10,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{run$\_$back}}=&\text{tolerance}(-\text{speed},(10,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{flip}}=&\text{tolerance}(\text{angmomentum},(5,\infty)).\\ \mathcal{R}_{\text{Cheetah}\_\text{flip$\_$back}}=&\text{tolerance}(-\text{angmomentum},(5,\infty)).\end{split}$ (32)
    Obviously, in this situation, for all states, the optimal actions of Cheetah-run and Cheetah-run__\__back are opposite, and TDR here is also huge.
Algorithm Hypothesis Cartpole-balance Walker-stand&walk Cheetah-run&run__\__back
&balance__\__sparse &prostrate&flip &flip&flip__\__back
CURL $\Pi_{1}$ 994.5 $\pm$ 3.6 254.1 $\pm$ 9.2 229.7 $\pm$ 10.9
CURL (w/ r) $\Pi_{3}$ 987.3 $\pm$ 12.9 265.0 $\pm$ 4.7 236.1 $\pm$ 5.8
SAC+AE $\Pi_{1}$ 992.5 $\pm$ 2.6 256.9 $\pm$ 5.9 225.8 $\pm$ 10.1
SAC+AE (w/ r) $\Pi_{3}$ 988.5 $\pm$ 7.0 257.2 $\pm$ 8.5 239.9 $\pm$ 12.4
Table 4: Performance (mean $\pm$ std) in DMC of CURL, CURL (w/ r), SAC+AE, and SAC+AE (w/ r).
Tasks Acrobot-Cartpole-Pendulum Walker-Cheetah-Hopper Cheetah-run__\__mass Walker-walk__\__mass
Train&Test Train&Test Train Test Train Test
Dreamer 541.3 $\pm$ 4.0 299.4 $\pm$ 17.2 717.8 $\pm$ 27.9 711.7 $\pm$ 38.8 889.9 $\pm$ 114.1 903.3 $\pm$ 102.8
TAD 667.7 $\pm$ 6.4 554.7 $\pm$ 23.6 754.3 $\pm$ 22.9 738.2 $\pm$ 41.1 957.8 $\pm$ 35.1 963.1 $\pm$ 32.6
[…]
C.2 Details of All Task Distributions
Now we introduce the three task distributions in our experiments, which are designed based on existing tasks in DeepMind control suite for testing the generalization of trained agents.
  • Cheetah__\__speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Cheetah__\__run in DeepMind control suite, and hopes the Cheetah robot can run with the target speed interval $(\alpha-\beta,\alpha+\beta)$.
    $\begin{split}&\mathcal{R}_{\text{Cheetah}\_\text{speed}}(\alpha,\beta)=\text{tolerance}(\text{speed},(\alpha-\beta,\alpha+\beta)).\end{split}$ (33)
    We train the agents in tasks with parameters ($0.5,0.2$), ($1.5,0.2$), ($2.0,0.2$), ($3.0,0.2$) and test them in tasks with parameters ($1.0,0.2$), ($2,5,0.2$).
  • Pendulum__\__angle($\alpha,\beta$). This task distribution is designed in this paper with parameter $-1\leq\alpha\leq\beta\leq 1$, based on the task Pendulumh__\__swingup in DeepMind control suite, and hopes the Pendulum robot can swing up within the target angle interval $(\arccos\alpha,\arccos\beta)$.
    $\begin{split}\mathcal{R}_{\text{Pendulum}\_\text{angle}}(\alpha,\beta)=\text{tolerance}(\text{angle},(&\arccos\alpha,\\ &\arccos\beta)).\end{split}$ (34)
    Training tasks are with parameters ($-0.95,-0.9$), ($-0.85,-0.8$), ($-0.8,-0.75$), ($-0.7,-0.65$) and test tasks are with parameters ($-0.9,-0.85$), ($-0,75,-0.7$).
  • Walker__\__speed($\alpha,\beta$). This task distribution is designed in this paper with parameter $0\leq\beta\leq\alpha$, based on the task Walker__\__run in DeepMind control suite, and hopes the Walker robot can run within the target speed interval $(\alpha-\beta,\alpha+\beta)$.
    $\begin{split}&\mathcal{R}_{\text{Walker}\_\text{speed}}(\alpha,\beta)=\text{tolerance}(\text{speed},(\alpha-\beta,\alpha+\beta)).\end{split}$ (35)
    We train the agents in tasks with parameters ($0.5,0.2$), ($1.5,0.2$), ($2.0,0.2$), ($3.0,0.2$) and test in tasks with parameters ($1.0,0.2$), ($2,5,0.2$).
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. For each seed and each task setting, it will take around 3 days. As for the hyper-parameters, we follow previous works (Ha and Schmidhuber 2018; Hafner et al. 2019b, a) and select 2 as the action repeat for all experiments following (Hafner et al. 2019a).
C.3 Visualization Results for Task Generalization
Moreover, for each task sampled from the task distribution Cheetah__\__speed (here parameters $(3.0,0.2)$, $(2.0,0.2)$, $(1.5,0.2)$, $(0.5,0.2)$ are for training tasks and parameters $(2.5,0.2),(0.5,0.2)$ are for test tasks), we plot the speed of the agent as a function of the timestep in Fig. 6. As depicted, for each task, the agent trained by TAD will quickly improve its speed until reaching the target speed and then keep its speed since the speed determines whether it has met the task requirements via utilizing historical information. Consequently, TAD is aware of different tasks and can successfully generalize to unseen test tasks. We also provide videos of these trajectories in supplementary materials.
Figure 6: Visualization of the trained TAD agent in the task distribution of Cheetah__\__speed. We plot the speed of the agent as a function of timesteps in all tasks.
Appendix D Experimental Details for MuJoCo
We here introduce state-based control tasks, including Half-CheetahFwd-Back, Half-Cheetah-Vel, and Humanoid-Direc-2D, in detail, following the setting of previous meta RL works (Finn, Abbeel, and Levine 2017; Rakelly et al. 2019).
  • Half-Cheetah-Fwd-Back. This task distribution includes two tasks: moving forward and moving backward.
  • Half-Cheetah-Vel. This task distribution hopes the agent to move forward and achieve the target velocity. There are 100 training tasks and 30 test tasks for experiments.
  • Humanoid-Direc-2D. This task distribution hopes the agent to move in the target direction. There are 100 training tasks and 30 test tasks for experiments.
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. For each seed and each task setting, it will take around 1 day. We select 1 as the action repeat for all experiments following.
$\beta$ 0.2 0.15 0.1
Train Test Train Test Train Test
Dreamer 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 247.5 $\pm$ 1.0 0.0 $\pm$ 0.0 175.8 $\pm$ 55.8 13.5 $\pm$ 13.5
TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 927.6 $\pm$ 2.6 800.1 $\pm$ 121.4 608.5 $\pm$ 321.4 491.6 $\pm$ 389.4
Table 6: Average cumulative reward (mean $\pm$ one std) over different target region (smaller $\beta$ represents smaller target region and more sparse return) of the best policy trained by Dreamer and TAD in Cheetah__\__Speed. For each $\beta$, we train agents in the train tasks and evaluate them in both train and test environments. Numbers greater than 95 percent of the best performance for each environment are bold.
SR 0.0 0.8 0.9
Train Test Train Test Train Test
Dreamer 250.2 $\pm$ 9.6 3.0 $\pm$ 2.2 237.4 $\pm$ 14.9 28.9 $\pm$ 14.6 168.7 $\pm$ 96.2 157.2 $\pm$ 188.4
TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 841.9 $\pm$ 154.9 546.1 $\pm$ 291.6 $7$$\pm$ 170.1 716.8 $\pm$ 136.5
Table 7: Average cumulative reward (mean $\pm$ one std) over different sparse rates of the best policy trained by Dreamer and TAD in Cheetah__\__Speed. For each sparse rate, we train agents in the train tasks and evaluate them in both train and test environments. Numbers greater than 95 percent of the best performance for each environment are bold.
Appendix E Ablation Study
E.1 Ablation Study on Reward Signals
In this part, we do ablation studies on reward signals for model-free methods including CURL and SAC+AE, which both belong to $\Pi_{1}$. We design CURL (w/ r) and SAC+AE (w/ r) by directly adding reward signals into the observation based on CURL and SAC+AE, respectively. As shown in Table. 4, the performance of CURL (w/ r) and SAC+AE (w/ r) is similar to CURL and SAC+AE. The major reason is that CURL (w/ r) and SAC+AE (w/ r)only utilize the observation and reward of the current timestep without historical information. Thus it is still difficult for them to distinguish different tasks.
E.2 Dynamic Generalization
As TAD utilizes all historical information to infer the environment, it can be directly applied to more general settings with different observations, dynamics, and/or actions. To evaluate the performance of TAD in these settings, we have conducted the following four experiments based on DMControl:
  • Acrobot-Cartpole-Pendulum: includes 7 tasks of artpole-balance, cartpole-balance__\__sparse, cartpole-swingup, cartpole-swingup__\__sparse, acrobot-swingup, acrobot-swingup__\__sparse, and pendulum-swingup. All these tasks aim to control a rod-shaped robot, while they own different embodiments, dynamics, and observations.
  • Walker-Cheetah-Hopper: includes 6 tasks of walker-prostrate, walker-stand, walker-walk, cheetah-run, hopper-stand, and hopper-hop. The tasks own different embodiments, dynamics, actions, and observations.
  • Cheetah-run-mass($m$): This task distribution is based on the task Cheetah-run and the mass of the robot is $m$ times that of the standard task. Thus different tasks own different dynamics. We train the agents in tasks with $m=0.6,1.0,1.2,1.6$ and test them in tasks with parameters $m=0.8,1.4$.
  • Walker-walk-mass($m$): This task distribution is based on the task Walker-walk and the mass of the robot is $m$ times that of the standard task. Thus different tasks own different dynamics. We train the agents in tasks with $m=0.6,1.0,1.2,1.6$ and test them in tasks with parameters $m=0.8,1.4$.
Then we test Dreamer and TAD in these four settings and report the results. TAD can achieve much greater performance and better convergence compared to Dreamer, as it can better infer the current task. This experiment indicates TAD’s potential in further handling dynamic generalization and even cross-embodiment tasks.
E.3 Sparse Reward
In this part, we will evaluate TAD in more challenging settings with sparse rewards. First, we evaluate Dreamer and TAD in Cheetah__\__speed with different $\beta$, which identifies the region of target speeds. With smaller $\beta$, the reward signals are more sparse since the target intervals are smaller. In the main experiment, we take $\beta=0.2$, and here we evaluate in $\beta=0.2,0.15,0.1$, of which the result is reported in Table 6. For each $\beta$, we take the training parameters ($0.5,\beta$), ($1.5,\beta$), ($2.0,\beta$), ($3.0,\beta$) and test them in tasks with parameters ($1.0,\beta$), ($2.5,\beta$). As shown in Table 6, with the decreasing of $\beta$, the performance of Dreamer and TAD decreases since reward signals are sparse so exploration here is much more difficult. However, our TAD still significantly outperforms Dreamer and shows strong generalization abilities, which shows that TAD can effectively utilize historical information, even sparse rewards.
Moreover, we design Cheetah__\__speed__\__sparse based on Cheetah__\__speed. In Cheetah__\__speed__\__sparse($n$), we make the reward function sparse, i.e., the output reward is the same as Cheetah__\__speed every $n$ timesteps (in step $n-1,2n-1,...$) and 0 otherwise, of which the sparse rate (SR) is $(n-1)/n$. We supplement experiments to evaluate the performance of Dreamer and TAD with $n=5$ (SR=0.8) and $n=10$ (SR=0.9). As shown in Table 7, with the increasing of SR, although the performance of TAD decreases since inferring task context from sparse reward is extremely difficult, TAD still shows strong performance in train tasks and generalizes well to unseen test tasks.
Appendix F Ethics Issues and Broader Impact
Designing agents that can generalize to unseen tasks is a major concern in reinforcement learning. This work focuses on task generalization in reinforcement learning and proposes a novel algorithm Reward Informed Dreamer. One of the potential negative impacts is that algorithms using deep neural networks, which lack interoperability and theoretical guarantee. If we hope to apply them in real-world applications, they may face security and robustness issues, and a possible way is to develop more explainable methods. There are no serious ethical issues as this is basic research. We hope our work can inspire more research on designing agents with stronger generalization abilities.